### v1

# Curriculum Approximate Unlearning for Session-based Recommendation

Abstract.
Approximate unlearning for session-based recommendation refers to eliminating the influence of specific training samples from the recommender without retraining of (sub-)models. Gradient ascent (GA) is a representative method to conduct approximate unlearning. However, there still exist dual challenges to apply GA for session-based recommendation. On the one hand, naive applying of GA could lead to degradation of recommendation performance. On the other hand, existing studies fail to consider the ordering of unlearning samples when simultaneously processing multiple unlearning requests, leading to sub-optimal recommendation performance and unlearning effect.
To address the above challenges, we introduce CAU, a curriculum approximate unlearning framework tailored to session-based recommendation. CAU handles the unlearning task with a GA term on unlearning samples. Specifically, to address the first challenge, CAU formulates the overall optimization task as a multi-objective optimization problem, where the GA term for unlearning samples is combined with retaining terms for preserving performance. The multi-objective optimization problem is solved through seeking the Pareto-Optimal solution, which achieves effective unlearning with trivial sacrifice on recommendation performance. To tackle the second challenge, CAU adopts a curriculum-based sequence to conduct unlearning on batches of unlearning samples. The key motivation is to perform unlearning from easy samples to harder ones. To this end, CAU first introduces two metrics to measure the unlearning difficulty, including gradient unlearning difficulty and embedding unlearning difficulty. Then, two strategies, hard-sampling and soft-sampling, are proposed to select unlearning samples according to difficulty scores. To verify the effectiveness of the proposed method, we conduct intensive experiments on three benchmark datasets and ablation studies show all the components and strategies of CAU contribute significantly to its superior effectiveness.
Unlearning, Session-based Recommendation, Privacy Protection, Recommender System, Information Security
submissionid: 452ccs: Information systems Recommender systemsccs: Security and privacy Web application security
1. Introduction
Session-based recommendation has shown effectiveness in predicting users’ future interests from historical interactions. Despite memorizing sequential patterns, the ability to eliminate the influence of specific training samples, a.k.a. recommendation unlearning (RU), is also crucially important. Several data protection regulations, such as the EU General Data Protection Regulation (GDPR), the California Consumer Privacy Act (CCPA), and the China Personal Information Protection Law (PIPL), have been proposed to ensure individuals’ rights to remove personal data from trained recommendation models (Mantelero, 2013; Baik, 2020). Beyond legal perspective, RU can further serve practical purposes, such as removing outdated user preferences (Liu et al., 2022; Matuszyk et al., 2015) or eliminating the impact of noisy training samples (Shaik et al., 2023).
Recommendation unlearning. Previous studies on RU can be categorized into exact recommendation unlearning and approximate recommendation unlearning (Li et al., 2024).
Exact recommendation unlearning attempts to completely eradicate the impact of the unlearning data by physically removing the unlearning samples and then retraining the model (Nguyen et al., 2022). To improve the unlearning efficiency, existing exact RU methods usually involve data partitioning and train sub-models on each data shard (Chen et al., 2022a, b; Bourtoule et al., 2019). The final results come from the aggregation of sub-models. However, when a large number of unlearning requests occur concurrently, the efficiency of exact RU could also deteriorate substantially, since the unlearning data are likely distributed across multiple shards, necessitating the retraining of several sub-models, or in the worst case, the entire model. Besides, due to prevalent item correlations, such exact RU methods may fail to achieve the exact unlearning effect (Xin et al., 2024). Even if the unlearning interaction has been removed from the data, it can still be inferred from correlated interactions due to the collaborative signals.
Approximate recommendation unlearning offers a more efficient alternative by enabling rapid forgetting and lowering the overall computational cost in contrast. Approximate recommendation unlearning denotes the process of efficiently removing the influence of specific training samples from a recommendation model without retraining the entire model or sub-models from scratch (Li et al., 2024). Gradient ascent (GA) is one of the most representative methods to conduct approximate recommendation unlearning. As deep learning models are typically trained using gradient descent-based optimization, an intuitive approach for unlearning is to get the opposite gradient for unlearning samples, and then add the gradient back to increase the loss function. GA mitigates the influence of unlearning samples from a parametric perspective, achieving the approximate unlearning effect. Yao et al. (2024) proposed to utilize GA-based methods to conduct unlearning for large language models.
Existing challenges. Despite its success, directly applying GA to conduct unlearning for session-based recommendation still poses non-trivial challenges.
i). Simply applying GA for session-based recommendation could result in significant degradation of recommendation performance. Since the objective function of GA is an inverted version of gradient descent, the objective function would diverge without tailored strategies, leading to failure of the entire training. Besides, due to the sequential correlations in session-based recommendation, conducting GA in one session could negatively impact the accuracy of other sessions. Thus, achieving a proper trade-off between unlearning and retaining through GA presents a key research challenge.
ii). Existing research fails to consider the sequence of unlearning samples when handling a batch of unlearning requests simultaneously. Curriculum learning has shown effectiveness in improving the learning process through learning from easy samples to hard ones. However, the effect of unlearning orders remains underexplored for session-based recommendation when handling a batch of unlearning requests. Intuitively, due to sequential correlations between interactions, the ordering of unlearning instances could significantly impact the unlearning outcome.
The proposed method. In this paper, we propose curriculum approximate unlearning (CAU), an approximate unlearning framework tailored to session-based recommendation. Concretely, the unlearning task is formulated as a GA term on unlearning samples. To address the first challenge, CAU defines the overall optimization task as a multi-objective optimization problem, where the GA term for unlearning samples is combined with retaining terms which aim to preserve the recommendation performance on remaining samples. Through seeking the Pareto-Optimal solution, CAU achieves effective unlearning with trivial sacrifice.
To address the second challenge, CAU uses a curriculum-based sequence to conduct unlearning on batches of unlearning samples. More precisely, we first propose two metrics to quantify the unlearning difficulty, including the gradient unlearning difficulty and embedding unlearning difficulty. The gradient unlearning difficulty reflects the consistency between the GA term gradient and the retaining term gradient, while the embedding unlearning difficulty captures semantic proximity between session states and unlearning items. Then, CAU introduces curriculum-based strategies to conduct unlearning according to the difficulty scores. Specifically, CAU introduces two unlearning strategies: (1) hard-sampling: a deterministic curriculum-based approach that strictly follows the unlearning difficulty ranking to select samples in every unlearning epoch, and (2) soft-sampling: a stochastic sampling strategy that selects easier samples with higher probability in early steps and gradually shifts towards harder ones in later steps.
To verify the effectiveness of the proposed method, CAU is implemented on three representative session-based recommendation models, including GRU4Rec, SASRec, and BERT4Rec. Extensive experiments on three benchmark datasets are conducted, and experimental results show the effectiveness of the proposed method.
Contributions. To summarize, the main contributions lie in:
  • To the best of our knowledge, CAU is the first attempt that conducts approximate unlearning for session-based recommendation through curriculum-based multi-task optimization. Pareto-Optimization is utilized to achieve the trade-off between unlearning and retaining recommendation performance.
  • We propose gradient unlearning difficulty and embedding unlearning difficulty to quantify unlearning difficulty in the context of session-based recommendation.
  • We propose two curriculum learning strategies to guide the unlearning process, starting with easier samples and progressively moving toward more difficult ones.
  • We conduct extensive experiments on three state-of-the-art session-based recommendation models and three benchmark datasets to show that CAU achieves efficient and effective unlearning while retaining high recommendation performance.
2. Related work
2.1. Session-based recommendation
Session-based recommendation aims to predict a user’s next action based on the current session. Early methods rely on session-based nearest-neighbor techniques, which recommend items frequently co-occurring with the current session  (Hariri et al., 2012, 2015; Bell and Koren, 2007). These methods are simple and interpretable but struggle to capture complex sequential patterns. With the rise of deep learning, RNN-based models  (Donkers et al., 2017; Hidasi and Karatzoglou, 2018; Mu, 2018) were proposed to model session sequences, achieving better performance by capturing temporal dependencies. Later works explored self-attention and Transformer-based architectures to model long-range dependencies and complex item interactions within sessions  (Kang and McAuley, 2018; Sun et al., 2019; Xu et al., 2019; Fang, 2021). Recent research further incorporates graph neural networks (GNNs) to represent sessions as graphs  (Wu et al., 2019; Wang et al., 2019; Qiu et al., 2019), capturing intricate item transition patterns and higher-order connectivity. Beyond model architecture, there are efforts on improving training and personalization strategies, including contrastive learning  (Liu et al., 2021; Xie et al., 2022), and meta-learning  (Song et al., 2021; Pan et al., 2022) for session-based recommendation. These approaches aim to better exploit session context, mitigate data sparsity, and improve generalization to unseen sessions.
2.2. Machine unlearning
Machine unlearning aims to remove the influence of certain training data from a given model. In traditional tasks like classification, previous studies have explored concepts related to machine unlearning, such as data deletion and selective forgetting  (Ginart et al., 2019; Guo et al., 2023; Izzo et al., 2021; Shibata et al., 2021). And current machine unlearning methods generally fall into two main categories: exact unlearning and approximate unlearning.
Exact unlearning methods typically rely on reversible training procedures or special model structures  (Cao and Yang, 2015; Bourtoule et al., 2019; Chen et al., 2022b), enabling the model to reconstruct its state as if the target samples were never seen. In contrast, approximate unlearning methods focus on efficiently reducing the residual impact of unlearning data, often by fine-tuning the model parameters with negative gradients  (Golatkar et al., 2020; Sekhari et al., 2021) or applying targeted weight perturbations  (Graves et al., 2021; Guo et al., 2019). While approximate methods cannot guarantee perfect removal, they offer significant advantages in scalability and computational efficiency  (Nguyen et al., 2022).
2.3. Recommendation unlearning
Unlearning in the recommendation scenario tends to attract more research attention. Recommendation unlearning can not only help to protect user privacy but also improve recommendation models through eliminating the effect of noisy data and misleading information  (Shaik et al., 2023).  Liu et al. (2022) and  Xu et al. (2023) proposed to use fine-tuning and the alternative least square algorithm for unlearning acceleration.  Chen et al. (2022a) and  Li et al. (2022) extended the ideas of the SISA algorithm for collaborative filtering. And  Hu et al. (2025) applied the similar idea to LLM-based recommendation unlearning.  Xin et al. (2024) proposed an exact unlearning framework for session-based recommendation. Additionally, there are also other techniques like matrix correction  (Liu et al., 2023), influence function  (Zhang et al., 2024) and label-flipping  (Alshehri and Zhang, 2023). However, none of the existing approximate unlearning methods is tailored for session-based recommendation. Previous works also do not take into account the difficulty of the unlearning samples, which may lead to suboptimal performance. In this work, we propose two methods to estimate the difficulty of unlearning samples and introduce two corresponding sampling strategies to implement a curriculum learning scheme, thereby further improving the effectiveness of the unlearning process.
3. Task definition
In this section, we describe session-based recommendation and then formulate the task of unlearning in this context.
3.1. Session-based recommendation
Let $V$ be the set of items, $S=\{v_{1},...,v_{t},...,v_{n}\}$ be the user-item interaction sequence in a session, where $v_{t}\in V$ is the item interacted at time step $t$ in the session. The dataset $D$ of user-item interaction sessions is defined as $D=\{S_{1},...,S_{|D|}\}$. Given a user-item interaction sequence at time step $t$, the learning objective of session-based recommendation is to predict the next item $v_{t+1}$ that the user is likely to interact with, which can be formalized as:
(1) $\mathbf{p}_{t+1}=\mathcal{M}_{\theta}(v|S,D)$
where $\mathbf{p}_{t+1}$ denotes the predicted probability distribution over candidate items at time step $t+1$, $\mathcal{M}_{\theta}$ represents the recommendation model parameterized by $\theta$, e.g., GRU4Rec (Hidasi et al., 2015) and SASRec (Kang and McAuley, 2018). Typically, the recommendation model $\mathcal{M}_{\theta}$ can be trained by optimizing a loss function $\mathcal{L}_{rec}$ based on the predicted probability distribution $\mathbf{p}_{t+1}$, such as the cross-entropy loss or the pair-wise ranking loss. The training process can be formalized as:
(2) $\theta^{rec}=\arg\min_{\theta}\mathcal{L}_{rec}(\theta^{inital};D)$
where $\theta^{initial}$ denotes the initial model parameters, and $\theta^{rec}$ represents the optimized parameters obtained by minimizing the recommendation loss $\mathcal{L}_{rec}$ over the dataset $D$.
Figure 1. Overview of the proposed CAU framework. CAU defines the unlearning objective $\mathcal{L}_{unlearn}$ and remaining terms $\mathcal{L}_{normal}$ and $\mathcal{L}_{KL}$. Then CAU introduces two metrics to measure the unlearning difficulty, including gradient difficulty and embedding difficulty. Two sampling methods are proposed to sample batches from easy to hard, according to the unlearning difficulty metrics. Finally, a Pareto-Optimality solver is used to perform multi-objective learning to obtain the unlearned model.
3.2. Unlearning
To address privacy concerns or enhance recommendation utility, recommender systems need to support unlearning requests to remove the impact of specific samples. For example, users might request to eliminate traces of interactions with sensitive items to preserve their privacy. Existing methods include exact recommendation unlearning and approximate recommendation unlearning.
Exact recommendation unlearning targets on completely eradicating the impact of the data to be forgotten as if they never occurred in the training process. Let $D_{f}\subset D$ denotes the data to be unlearned, the naive approach to achieve exact unlearning is to retrain the model from scratch on the dataset without the unlearning data, i.e., $D-D_{f}$. This process can be defined as:
(3) $\theta^{exa}=\arg\min_{\theta}\mathcal{L}_{rec}(\theta^{inital};D-D_{f})$
where $\theta^{exa}$ represents the parameters of exact unlearned model obtained by minimizing the recommendation loss $\mathcal{L}_{rec}$ over $D-D_{f}$.
In contrast to exact unlearning, approximate recommendation unlearning offers a more computationally efficient alternative that aims to emulate the effect of exact unlearning, while preserving model utility and avoiding retraining. Instead of retraining the model from scratch, approximate unlearning optimizes the model parameters directly in response to unlearning requests. Common methods for approximate unlearning include gradient ascent on unlearning samples to reverse their influence (Graves et al., 2021; Sekhari et al., 2021) or filtering the model’s outputs (Baumhauer et al., 2022). Approximate recommendation unlearning can be defined as:
(4) $\theta^{app}=\arg\min_{\theta}\mathcal{L}_{app}(\theta^{rec};D_{f})$
where $\theta^{app}$ denotes the updated model parameters obtained by optimizing a specially designed approximate unlearning loss $\mathcal{L}_{app}$, starting from the previous recommendation model parameters $\theta^{rec}$.
4. Method
In this section, we detail the operational mechanism of CAU, which primarily consists of multi-objective learning and curriculum-based sample selection.
The framework is illustrated in Figure 1. Specifically, to maintain recommendation performance after unlearning, CAU formulates the task as a multi-objective learning problem. CAU incorporates a learning objective that focuses on preserving recommendation performance, alongside the unlearning objective for the unlearning data $D_{f}$. We then delve into the model’s gradient and embedding space and introduce two metrics to measure unlearning difficulty, respectively. Based on these difficulty metrics, CAU proposes two sampling methods—hard-sampling and soft-sampling—to construct unlearning batches, which facilitates a curriculum strategy to perform unlearning from easy to hard. Finally, a Pareto-Optimality solver is utilized to achieve multi-objective optimization to obtain the unlearned model parameters.
4.1. Multi-objective unlearning
While effective unlearning is important, maintaining recommendation performance is equally crucial. This necessitates a careful balance between the unlearning objective and the recommendation utility, so we formulate the approximate unlearning as a multi-task optimization problem. The key idea is to jointly optimize the recommendation utility and unlearning effectiveness through training a composite objective function with three task-specific losses: (1) the unlearning loss $\mathcal{L}_{unlearn}$ that encourages forgetting specific samples, (2) the retention loss $\mathcal{L}_{normal}$ to preserve recommendation performance, and (3) the consistency loss $\mathcal{L}_{KL}$ that regularizes deviations from the previous model.
Next, we will introduce the three types of loss functions in detail and describe how they are weighted during optimization.
Unlearn Loss. Unlearning can be regarded as the inverse process of forward learning, so we apply gradient ascent to explicitly remove the influence of the unlearning sample, which intuitively corresponds to adding back the gradient of the target data that was previously subtracted. For the target item $v_{t}$ to be unlearned, the unlearning loss is formulated as:
(5) $\mathcal{L}_{unlearn}=log(P_{\theta}(v_{t}|v_{1:t-1}))$
This formulation effectively pushes the model away from favoring the unlearning item, by reducing its predicted probability in the output distribution.
Normal Loss. To ensure that the unlearned model retains its general predictive capability, we include a standard cross-entropy loss based on the next item $v_{t+1}$, given the prefix session $\{v_{1},...,v_{t-1}\}$:
(6) $\mathcal{L}_{normal}=-log(P_{\theta}(v_{t+1}|v_{1:t-1}))$
Since unlearning is performed to remove an item from a session, it is essential to ensure that the model maintains continuity on remaining items in the session. $\mathcal{L}_{normal}$ guides the model to remain functional in its primary task, i.e., predicting plausible future items based on user behavior, despite forgetting specific items.
KL Loss. While removing unwanted information, CAU aims to minimize catastrophic drift from the previous model to better maintain the model’s recommendation utility. Therefore, we include a KL divergence loss to align the output distribution of the unlearned model $P_{\theta}$ with the previous model $P_{\theta^{rec}}$:
(7) $\mathcal{L}_{KL}=KL(P_{\theta^{rec}}(\cdot|v_{1:t-1})||P_{\theta}(\cdot|v_{1:t-1}))$
This ensures that for remaining samples, the behavior of the new model does not deviate significantly from the previous model, improving stability and recommendation consistency.
Finally, the loss function for multi-objective unlearning is formulated as follows:
(8) $\mathcal{L}_{app}=\alpha_{1}\mathcal{L}_{unlearn}+\alpha_{2}\mathcal{L}_{normal}+\alpha_{3}\mathcal{L}_{KL}$
4.2. Pareto-Optimality solver
A crucial challenge in multi-task optimization is balancing the importance of each objective. Manually tuning the weights $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ is suboptimal and sensitive to data distribution and model scale. Furthermore, maintaining fixed hyperparameters throughout the entire training process can hinder the model’s ability to adapt to changes and negatively impact performance. To effectively solve this problem, we employ a Pareto-Otimality solver, which can reliably find a stable solution that balances the trade-offs between conflicting objectives. This approach ensures that the model simultaneously achieves effective unlearning while maintaining overall recommendation quality. Given a model parameterized by $\theta$ and a series of learning objective functions $\mathcal{L}_{unlearn},\mathcal{L}_{normal},\mathcal{L}_{KL}$, Pareto-Optimality is a state in which it is impossible to enhance one objective without compromising another. Formally, the Pareto-Optimality can be defined as:
Definition 4.1.
A solution $\theta$ dominates $\bar{\theta}$ if $\mathcal{L}_{k}(\theta)\leq\mathcal{L}_{k}(\bar{\theta})$ for all learning objective functions $\mathcal{L}_{k}$.
Definition 4.2.
A solution $\theta^{*}$ is called Pareto-Optimal if there exists no solution $\theta$ that dominates $\theta^{*}$.
Given the training objective $\mathcal{L}_{app}$ in Eq.(8), CAU aims to find $\alpha_{i}$ so that the solution is Pareto-Optimal. Therefore, we employ the Multiple Gradient Descent Algorithm (MGDA) to find a solution that balances the competing objectives. MGDA  (Sener and Koltun, 2018) seeks a convex combination of gradients from all objectives such that their weighted sum forms a common descent direction.
Specifically, at each training step, we compute the gradients of each individual task loss w.r.t. the model parameters $\theta$:
(9) $\mathbf{g}_{1}=\nabla_{\theta}\mathcal{L}_{unlearn},\quad\mathbf{g}_{2}=\nabla_{\theta}\mathcal{L}_{normal},\quad\mathbf{g}_{3}=\nabla_{\theta}\mathcal{L}_{KL}.$
Formally, MGDA uses the Karush-Kuhn-Tucker (KKT) conditions (Bertsekas, 1997) to describe constraints of finding $\alpha_{1},\alpha_{2},\alpha_{3}$:
(10) $\min_{\alpha_{1},\alpha_{2},\alpha_{3}}\left\|\sum_{i=1}^{3}\alpha_{i}\mathbf{g}_{i}\right\|^{2}_{2}\quad\text{subject to}\quad\sum_{i=1}^{3}\alpha_{i}=1,\ \alpha_{i}\geq 0$
Employing MGDA solver to Eq.(10), CAU obtains optimal task weights $\alpha_{i}$ for each learning objective, ensuring that the training progress is in a direction which improves all tasks without favoring one while sacrificing others.
After computing $\alpha_{i}$ for all tasks, we utilize the loss weighted by $\alpha_{i}$ to update the model parameters, thus identifying an optimization direction that preserves the performance of recommendation while improving the effectiveness of unlearning.
4.3. Curriculum unlearning
To further enhance effectiveness of approximate unlearning when handling multiple requests simultaneously, we propose a curriculum-based unlearning strategy that prioritizes forgetting easier samples first, progressively moving toward harder ones. The underlying assumption is that removing low-impact easier samples earlier helps stabilize training and facilitates optimization. The curriculum unlearning framework comprises two key components: (1) difficulty estimation and (2) sample scheduling.
4.3.1. Difficulty estimation
In this study, we consider two perspectives for measuring unlearning difficulty, including gradient unlearning difficulty and embedding unlearning difficulty.
Gradient unlearning difficulty. This metric evaluates how difficult it is to update the model in a direction that jointly improves (or at least does not degrade) all individual tasks. Since we adopt MGDA for multi-objective optimization, large conflicts between the gradients of individual objectives would hinder the search for a Pareto-Optimal solution.
To quantify this, we compute the cosine similarity between the unlearning gradient and the sum of the gradients of the normal and KL losses. Specifically, for an unlearning sample $x$, the gradient unlearning difficulty is defined as:
(11) $Dif_{g}(x)=-\cos\left(\nabla_{\theta}\mathcal{L}_{unlearn}(x),\ \nabla_{\theta}\mathcal{L}_{normal}(x)+\nabla_{\theta}\mathcal{L}_{KL}(x)\right)$
A higher cosine similarity implies better alignment among the objectives, meaning the sample is easier to unlearn. Therefore, we sort the samples by increasing cosine similarity and prioritize those with smaller gradient conflicts in the beginning.
Gradient unlearning difficulty measures the difficulty from an external perspective, i.e., measuring the degree of conflicts between unlearning and retaining.
Embedding unlearning difficulty. This metric is based on the intuition that samples which the model strongly favors are harder to be forgotten. Specifically, we compute the dot product between the session state representation and the embedding of the item to be forgotten. Given a session state representation $\mathbf{e}_{session}$ and the embedding of the target item $x$ to be forgotten $\mathbf{e}_{x}$, CAU computes:
(12) $Dif_{e}(x)=\langle\mathbf{e}_{session},\ \mathbf{e}_{x}\rangle$
Here, $\mathbf{e}_{session}$ is the encoded state representation of the session prefix (e.g., from an RNN or self-attention network), and $\mathbf{e}_{x}$ is the embedding of the target item $x$ to be forgotten. $\langle\rangle$ denotes dot product. A higher $Dif_{e}$ implies the model is more confident in predicting the unlearning target $x$ , making it harder to be removed. Hence, we treat samples with lower $Dif_{e}$ as easier ones and schedule them earlier in the unlearning process.
Input: unlearning data $D_{f}$, training epochs $E$, recommender $\mathcal{M}$ with parameter $\theta^{rec}$
Output: unlearned model $\mathcal{M}_{u}$ with parameter $\theta^{app}$
1
[…]
3 for $epoch\leftarrow 1$ to $E$ do
4   calculate $Dif_{g}$ or $Dif_{e}$ according to Eq.(11) or Eq.(12);
5 $D_{f\text{-sorted}}\leftarrow\text{sort }D_{f}\text{ ascendingly }\text{by $Dif_{g}$ or $Dif_{e}$}$;
6   sample batchbatch $B$ sequentially from $D_{f\text{-sorted}}$;
7 foreach batch $B$ do
8    $\alpha_{1},\alpha_{2},\alpha_{3}\leftarrow$ calculate weights by Eq.(10);
9    $\mathcal{L}_{app}\leftarrow$ calculate loss using Eq.(8);
10    $\theta^{app}\leftarrow$ update model parameter according to $\mathcal{L}_{app}$;
11
12   end foreach
13
14 end for
return $\mathcal{M}_{u}$ with $\theta^{app}$
Algorithm 1 CAU training with hard-sampling
Embedding unlearning difficulty measures the difficulty from an internal perspective, i.e., whether the recommender itself tends to recommend the target unlearning item $x$.
In practice, both difficulty metrics can be used to guide the following curriculum sampling for unlearning.
4.3.2. Sample scheduling
CAU introduces two strategies to achieve the unlearning curriculum: hard-sampling and soft-sampling.
Hard-sampling. In the hard-sampling strategy, the training process for each epoch deterministically adheres to the easy-to-hard curriculum schedule. Specifically, all samples in the unlearning dataset are firstly ranked according to their difficulty scores. The training then proceeds sequentially following the order, beginning with samples deemed easier to be forgotten and gradually advancing towards more challenging samples.
For example, when employing the gradient unlearning difficulty metric, the initial phase of each epoch concentrates on samples with high gradient alignment (i.e., lower $Dif_{g}$), indicating that their gradients align well and thus are easier to be optimized. As training progresses, more challenging samples with higher $Dif_{g}$ and more conflicting gradients are incrementally introduced into the training batches. Such curriculum progression reduces abrupt gradient conflicts and helps maintain a stable optimization trajectory.
The detailed procedure of the hard-sampling strategy is formalized and presented in Algorithm 1.
Soft-sampling. To introduce stochasticity into the unlearning process and mitigate the risk of overfitting to a fixed curriculum, CAU proposes a soft-sampling mechanism where samples are drawn stochastically at each step based on their difficulty scores. Unlike hard-sampling which deterministically follows a strict curriculum order, soft-sampling assigns a dynamic probability distribution over the samples, enabling more flexible and robust training.
Formally, the sampling probability for each sample $x$ at time step $t$ is defined as follows:
(13) $p_{x}(t)=\frac{(\exp((2t-1)\cdot(Dif(x)-\bar{Dif})))^{\tau}}{\sum_{j}(\exp((2t-1)\cdot(Dif(j)-\bar{Dif})))^{\tau}}$
Input: unlearning data $D_{f}$, total training steps $T$, recommender $\mathcal{M}$ with parameter $\theta^{rec}$
Output: unlearned model $\mathcal{M}_{u}$ with parameter $\theta^{app}$
1
[…]
3
4for $step\leftarrow 1$ to $T$ do
5 $t\leftarrow step/T$;
6 foreach sample $x$ in $D_{f}$ do
7      calculate DifDif according to Eq.(11) or Eq.(12);
8    $p_{x}(t)\leftarrow$ calculate the sampling probability according to Eq.(13);
9
10   end foreach
11
12 $B\leftarrow$ sample a batch from $D_{f}$ according to $p_{x}(t)$;
13
14 $\alpha_{1},\alpha_{2},\alpha_{3}\leftarrow$ calculate weights by Eq.(10);
15 $\mathcal{L}_{app}\leftarrow$ calculate loss using Eq.(8);
16 $\theta^{app}\leftarrow$ update model parameter according to $\mathcal{L}_{app}$;
17
18 end for
return $\mathcal{M}_{u}$ with $\theta^{app}$
Algorithm 2 CAU training with soft-sampling
where $Dif$ denotes the difficulty scores (i.e., either $Dif_{g}$ or $Dif_{e}$), and $\bar{Dif}$ is the mean difficulty score of all unlearning samples. The time variable $t$ is normalized to the interval $[0,1]$ based on the current training progress, calculated as t=current_step/total_stept=current\_step/total\_step. $\tau$ is the temperature coefficient.
Soft-sampling ensures that at the early stages of training ($t\approx 0$), the exponent (2t1)(2t-1) is close to 1-1, which inversely weights the difficulty score, making easier samples with lower difficulty scores more likely to be selected. As training progresses ($t\approx 1$), the exponent (2t1)(2t-1) approaches +1+1, shifting the sampling preference toward harder samples. This step-dependent adjustment of the sampling distribution can be viewed as a form of curriculum learning with soft transitions, allowing the model to gradually focus on increasingly difficult samples without abrupt switches. Algorithm 2 illustrates the overall procedure of soft-sampling.
5. Experiments
5.1. Experimental settings
5.1.1. Datasets and pre-processing
We conduct experiments on three public accessible datasets: MovieLen-1M [^1] https://grouplens.org/datasets/movielens/1m/, Amazon Beauty and Games [^2] https://jmcauley.ucsd.edu/data/amazon/. These datasets vary in scale, domain, and sparsity.
Following the data pre-processing procedure of (Sun et al., 2019), we filter out users and items with fewer than 5 interactions and sort user interactions chronologically to form sequences. For each dataset, we adopt cross-validation to evaluate the performance, and we randomly split sessions into training set, validation set, and test set with a ratio of 8:1:1.
5.1.2. Evaluation protocols
The recommendation performance is measured with two metrics: $Recall@k$ and Normalized Discounted Cumulative Gain@k@k ($NDCG@k$). $Recall@k$ measures how many ground-truth items are included in the top-$k$ positions of the recommendation list. $NDCG@k$ is a rank-sensitive metric that assigns higher weights to top positions in the recommended list.
To evaluate unlearning effectiveness, we adopt the $Hit_{u}@k$ metric introduced in (Xin et al., 2024). It is used to measure the unlearning effect through assessing whether the unlearned item reappears in the top-$k$ recommendation list generated by the unlearned model, based on the remaining interactions within the session. Lower $Hit_{u}@k$ denotes better unlearning effectiveness.
Given that recommendation performance and unlearning effectiveness follow distinct numerical patterns, such as higher $NDCG@k$ and $Recall@k$ indicate better recommendation performance, while lower $Hit_{u}@k$ signifies better unlearning effectiveness. We introduce a new metric $U_{\beta}$ to facilitate a comprehensive evaluation of both aspects. Specifically, $U_{\beta}$ integrates $Recall@k$ and $Hit_{u}@k$ as:
(14) $U_{\beta}=(1+\beta^{2})\cdot\frac{Recall\times(1-Hit_{u})}{\beta^{2}\times Recall+(1-Hit_{u})}.$
The hyperparameter $\beta$ controls the relative emphasis between recommendation performance and unlearning effectiveness. This formulation enables a comprehensive evaluation that reflects the trade-off between maintaining high recommendation quality and ensuring effective removal of specific items from the model’s output.
5.1.3. Baselines
Three representative session-based recommendation models are used as the backbone models.
  • GRU4Rec (Hidasi et al., 2015) utilizes gated recurrent units (GRU) to model user interaction sequences which is an RNN-based sequential recommendation model.
  • SASRec (Kang and McAuley, 2018) uses the left-to-right unidirectional Transformer (Vaswani et al., 2017) decoder for session-based recommendation.
  • BERT4Rec (Sun et al., 2019) employs bidirectional self-attention to model interaction sequences.
Table 1. Recommendation performance comparison after unlearning 10% of data. N is short for NDCG, R is short for Recall. Best results other than Original and Retrain are highlighted in bold.
Dataset / Method GRU4Rec SASRec BERT4Rec
Beauty N@10 N@20 R@10 R@20 N@10 N@20 R@10 R@20 N@10 N@20 R@10 R@20
Original 0.0317 0.0368 0.0533 0.0739 0.0264 0.0325 0.0513 0.0756 0.0270 0.0339 0.0511 0.0781
Retrain 0.0326 0.0375 0.0562 0.0756 0.0286 0.0350 0.0512 0.0770 0.0256 0.0314 0.0496 0.0728
SISA 0.0107 0.0135 0.0153 0.0262 0.0055 0.0082 0.0120 0.0225 0.0089 0.0179 0.0103 0.0232
SRU 0.0191 0.0226 0.0332 0.0455 0.0178 0.0223 0.0332 0.0511 0.0175 0.0172 0.0332 0.0409
CAU+gradient+hard 0.0272 0.0333 0.0484 0.0725 0.0271 0.0327 0.0527 0.0747 0.0245 0.0307 0.0460 0.0710
CAU+gradient+soft 0.0254 0.0308 0.0485 0.0700 0.0259 0.0320 0.0513 0.0758 0.0246 0.0302 0.0475 0.0695
CAU+embedding+hard 0.0289 0.0349 0.0529 0.0769 0.0267 0.0313 0.0536 0.0717 0.0245 0.0308 0.0465 0.0697
CAU+embedding+soft 0.0253 0.0304 0.0483 0.0686 0.0255 0.0308 0.0522 0.0732 0.0240 0.0300 0.0440 0.0678
Games N@10 N@20 R@10 R@20 N@10 N@20 R@10 R@20 N@10 N@20 R@10 R@20
Original 0.0569 0.0713 0.1049 0.1618 0.0461 0.0590 0.1004 0.1511 0.0565 0.0706 0.1096 0.1660
Retrain 0.0593 0.0703 0.0110 0.1538 0.0498 0.0640 0.1041 0.1607 0.0528 0.0666 0.1012 0.1557
SISA 0.0092 0.0115 0.0170 0.0261 0.0119 0.0149 0.0196 0.0318 0.0172 0.0189 0.0203 0.0515
SRU 0.0197 0.0226 0.0337 0.0485 0.0266 0.0239 0.0501 0.0569 0.0293 0.0361 0.0557 0.0834
CAU+gradient+hard 0.0553 0.0680 0.1066 0.1572 0.0489 0.0640 0.0940 0.1544 0.0500 0.0651 0.1001 0.1555
CAU+gradient+soft 0.0523 0.0656 0.1011 0.1544 0.0489 0.0627 0.0960 0.1513 0.0502 0.0645 0.0969 0.1548
CAU+embedding+hard 0.0552 0.0672 0.1053 0.1534 0.0493 0.0638 0.0967 0.1529 0.0508 0.0655 0.0997 0.1586
CAU+embedding+soft 0.0534 0.0651 0.1059 0.1524 0.0502 0.0633 0.0974 0.1522 0.0500 0.0635 0.0963 0.1534
ML-1M N@10 N@20 R@10 R@20 N@10 N@20 R@10 R@20 N@10 N@20 R@10 R@20
Original 0.1315 0.1521 0.2361 0.3171 0.1243 0.1482 0.2155 0.3314 0.1302 0.1538 0.2315 0.3261
Retrain 0.1269 0.1512 0.2241 0.3201 0.1325 0.1590 0.2454 0.3546 0.1293 0.1564 0.2300 0.3377
SISA 0.0212 0.0343 0.0391 0.0915 0.0185 0.0254 0.0317 0.0592 0.0172 0.0242 0.0203 0.0416
SRU 0.0747 0.0919 0.1255 0.1938 0.0430 0.0497 0.0660 0.0966 0.0412 0.0506 0.0611 0.0977
CAU+gradient+hard 0.1280 0.1485 0.2311 0.3096 0.1129 0.1328 0.2132 0.2928 0.1064 0.1286 0.2010 0.2943
CAU+gradient+soft 0.1329 0.1531 0.2351 0.3103 0.1018 0.1206 0.1910 0.2644 0.1080 0.1284 0.2063 0.2825
CAU+embedding+hard 0.1333 0.1514 0.2357 0.3075 0.1136 0.1347 0.2088 0.2918 0.1104 0.1309 0.2048 0.2906
CAU+embedding+soft 0.1318 0.1485 0.2326 0.3000 0.1018 0.1242 0.1926 0.2825 0.1156 0.1300 0.2075 0.2946
Each model is trained with the following model-agnostic unlearning frameworks:
  • Original: the original model without unlearning.
  • Retrain: retrain the whole model from scratch on the remaining dataset. It’s computationally expensive.
  • SISA (Bourtoule et al., 2019): a fundamental exact unlearning method that randomly splits the data and averages the outputs of the sub-model.
  • SRU (Xin et al., 2024): a newly proposed method which improves SISA by collaborative sharding and extra deletion.
  • CAU: the proposed method, comprising the following variants:
      gradient: use the gradient unlearning difficulty.
      embedding: use the embedding unlearning difficulty.
      hard: use the hard-sampling strategy.
      soft: use the soft-sampling stragety.
5.1.4. Implementation details
All methods are learned with the Adam optimizer (Kingma and Ba, 2014). The learning rate is set as 1e31e-3. The model input consists of the most recent 10 interacted items for Beauty, 20 for Games, and 50 for MovieLens-1M, with shorter sequences padded using a special padding token. The training batch size is set to 256, while the unlearning batch size is set to 128. Temperature coefficient $\tau$ is 2. The embedding size is fixed at 64. For the unlearning process in hard-sampling, the unlearning epoch is set to 200 for SASRec and BERT4Rec, and 100 for GRU4Rec. And for soft-sampling, the number of unlearning step is adjusted correspondingly. Each experiment is repeated 5 times, and the average performance is reported.
Table 2. Unlearning effectiveness and overall performance comparison of different methods on the three datasets. The U-score is computed by Recall@10 and $\text{Hit}_{u}@1$, with $\beta=10$ for Beauty and $\beta=3$ for Games and ML-1M. Boldface denotes the best score.
Dataset / Method GRU4Rec SASRec BERT4Rec
[…]
Original 0.6961 0.8950 0.2904 0.6355 0.8319 0.3437 0.2169 0.4285 0.6857
Retrain 0.1470 0.2685 0.7479 0.1520 0.2822 0.7347 0.1408 0.2853 0.7396
SISA 0.0648 0.1640 0.5857 0.0579 0.1569 0.5330 0.0489 0.1457 0.6273
SRU 0.0822 0.1918 0.7261 0.0828 0.1968 0.7259 0.0775 0.1740 0.7291
CAU+gradient+hard 0.1262 0.2007 0.7476 0.0973 0.1821 0.7784 0.1324 0.2648 0.7373
CAU+gradient+soft 0.1088 0.1681 0.7603 0.0823 0.1647 0.7862 0.1296 0.2564 0.7430
CAU+embedding+hard 0.1125 0.2094 0.7675 0.0954 0.1784 0.7817 0.1290 0.2672 0.7410
CAU+embedding+soft 0.1010 0.1793 0.7655 0.0848 0.1635 0.7865 0.1283 0.2632 0.7349
[…]
Original 0.4721 0.7669 0.3762 0.4061 0.7022 0.3982 0.2897 0.5765 0.4588
Retrain 0.2113 0.4278 0.4884 0.2109 0.4389 0.4759 0.1875 0.4129 0.4771
SISA 0.0632 0.1610 0.1463 0.0533 0.1760 0.1654 0.0524 0.1589 0.2255
SRU 0.0816 0.1912 0.2534 0.0691 0.1799 0.3374 0.0691 0.1620 0.3622
CAU+gradient+hard 0.1619 0.3196 0.4970 0.1521 0.3353 0.4704 0.1679 0.3596 0.4806
CAU+gradient+soft 0.1449 0.2765 0.4898 0.1410 0.3221 0.4787 0.1585 0.3507 0.4761
CAU+embedding+hard 0.1525 0.3161 0.4971 0.1423 0.3255 0.4799 0.1636 0.3562 0.4810
CAU+embedding+soft 0.1457 0.2867 0.5006 0.1402 0.3034 0.4823 0.1585 0.3477 0.4743
[…]
Original 0.3126 0.5983 0.5771 0.2774 0.6025 0.5850 0.2857 0.6066 0.5844
Retrain 0.2733 0.5694 0.5936 0.2650 0.5673 0.6127 0.2464 0.5404 0.6139
SISA 0.0212 0.0841 0.2876 0.0414 0.1636 0.2443 0.0207 0.0849 0.1708
SRU 0.0714 0.2059 0.5662 0.0538 0.2215 0.4055 0.0042 0.0798 0.3935
CAU+gradient+hard 0.1325 0.2987 0.6802 0.1366 0.2899 0.6616 0.0932 0.1967 0.6711
CAU+gradient+soft 0.1201 0.2650 0.6905 0.1139 0.2857 0.6497 0.0642 0.1905 0.6913
CAU+embedding+hard 0.1284 0.2547 0.6864 0.1366 0.3043 0.6573 0.1035 0.2298 0.6702
CAU+embedding+soft 0.0994 0.2526 0.6997 0.1118 0.2692 0.6525 0.0911 0.2257 0.6793
5.2. Recommendation performance
We list the recommendation performance comparison results in Table 1. We can observe that due to the removal of 10% of the training data, the recommendation performance declines across all baselines, which is an expected outcome. However, in the context of the unlearning task, our primary concern is the ability to preserve recommendation performance on the remaining samples rather than the unlearning effectiveness alone.
From the table, We observe that the recommendation performance of SISA and SRU drops markedly, possibly due to the partitioning of the dataset into disjoint shards, which substantially reduces correlations between sessions. Specifically, even with the incorporation of attention-based aggregation mechanisms in SRU, its effectiveness in mitigating performance degradation remains limited. Compared with Retrain, their performance drops by 63% and 32%, respectively. Such a drastic reduction is undesirable, as overly low recommendation performance would directly impair the user experience.
In contrast, CAU demonstrates significantly superior performance preservation, achieving results close to the Original model with a stable performance drop controlled within 8%. This advantage can be attributed to the introduction of the normal loss and KL loss during the unlearning model update, which suppresses excessive disturbance to the original representation distribution, thereby enabling effective unlearning of target samples while retaining the model’s recommendation capability as much as possible.
5.3. Unlearning effectiveness and efficiency
5.3.1. Unlearning effectiveness.
The results of unlearning performance are summarized in Table 2. The Original model represents the effect without unlearning, showing very high $Hit_{u}$ values. Retrain is the most ideal yet computationally expensive approach, but due to the abundant correlations and connections inherent in session-based recommendation, its $Hit_{u}$ value is not always the lowest. However, the recommendation performance of SISA and SRU drops significantly. In this case, even if the $Hit_{u}$ metric is low, it holds little practical significance. To enhance comparability among models, the U-score is employed to jointly measure both recommendation performance and unlearning effectiveness. In the overall comparison, Retrain achieves a relatively high U-score as it attains a well-balanced trade-off between recommendation performance and unlearning effectiveness. Although the Original model exhibits strong recommendation performance, it performs poorly in unlearning, resulting in a low U-score. Both SISA and SRU suffer from unstable and suboptimal U-scores due to their inability to maintain adequate recommendation quality. But CAU consistently achieves the best performance, maintaining recommendation accuracy close to Original while attaining a lower $Hit_{u}$ score than Retrain.
What’s more, we can observe that in most cases, soft-sampling is more effective than hard-sampling. For instance, on the ML-1M and BERT4Rec datasets using gradient unlearning difficulty, the U-score of soft-sampling is 0.6913 compared to 0.6711 for hard-sampling. Similarly, with embedding unlearning difficulty, soft-sampling consistently achieves higher scores than hard-sampling. This improvement is likely due to the increased robustness that soft-sampling provides throughout the unlearning procedure.
Table 3. Comparison of unlearning efficiency (minute [m]). Boldface denotes the best score.
Method Retrain SISA SRU CAU
GRU4Rec 30.72m 3.46m 5.45m 0.57m
SASRec 29.21m 4.54m 9.04m 2.24m
BERT4Rec 30.24m 4.64m 8.47m 2.24m
5.3.2. Unlearning efficiency.
Table 3 compares the unlearning time of the three models on the Games dataset. It can be observed that the Retrain method is undoubtedly the most time-consuming. For SRU and SISA, exact unlearning requires retraining sub-models or aggregation layers, resulting in substantial time overhead. In contrast, CAU achieves the best overall efficiency by training exclusively on the unlearning data while attaining satisfactory unlearning performance within no more than 200 epochs. In most cases, CAU achieves a speedup of approximately $13\times$ over Retrain and $3\times$ over SISA and SRU, which further demonstrates the efficiency of our approximate unlearning method.
Figure 2. Impact of the unlearning batch size.
Table 4. Ablation study on three datasets with SASRec as backbone model. Boldface denotes the best scores. w/o denotes without. N is short for NDCG, R is short for Recall.
[…]
Beauty w/o PS 0.0236 0.0503 0.0976 0.7727
w/o CU 0.0254 0.0499 0.1106 0.7624
w/o RL 0.0010 0.0027 0.0165 0.2127
CAU 0.0255 0.0522 0.0848 0.7865
Games w/o PS 0.0453 0.0845 0.0486 0.4695
w/o CU 0.0480 0.0933 0.1572 0.4673
w/o RL 0.0009 0.0027 0.0030 0.0256
CAU 0.0502 0.0974 0.1402 0.4823
ML-1M w/o PS 0.0769 0.1526 0.0269 0.6329
w/o CU 0.1021 0.1966 0.1491 0.6384
w/o RL 0.0012 0.0031 0.0041 0.0302
CAU 0.1129 0.2132 0.1366 0.6616
5.4. Ablation study
In this part, we conduct ablation study to analyze the functionality of the main components and strategies of CAU (i.e., Pareto-Optimality solver (PS), curriculum unlearning (CU), and recommendation loss (RL), where RL consists of both the normal loss and the KL loss). Table 4 shows the performance of CAU and its variants on all three datasets using SASRec as the backbone model. We introduce the variants and analyze the effects respectively.
(1) Effect of Pareto-Optimality solver: We removed the Pareto-Optimality solver component and assigned an equal weight to each task, which means setting all $\alpha$ values to 0.33. The experimental results demonstrate the effectiveness of the Pareto-Optimality solver: by formulating the entire unlearning process as a multi-task optimization problem, the Pareto-Optimality solver iteratively searches for gradient update directions that do not degrade any individual sub-task, thereby achieving a better balance between recommendation performance and unlearning effectiveness.
(2) Effect of curriculum unlearning: We remove the curriculum unlearning strategy and unlearn the data randomly in each epoch. This proves the effectiveness of our proposed method: guiding the model to unlearn data from easy to difficult based on their unlearning difficulty scores enables it to achieve a satisfactory forgetting level more quickly.
(3) Effect of recommendation loss: We remove both the normal loss and the KL loss, which means directly apply GA to unlearn data. From the results we can find that this will lead to catastrophic effects, as the model’s recommendation accuracy would almost completely collapse within fewer than five unlearning epochs, and the $Hit_{u}$ value would become meaningless. This observation highlights the critical role of the recommendation loss in preserving the model’s recommendation capability.
5.5. Hyperparameter study
5.5.1. Impact of unlearning batch size
In this part, we investigate the impact of batch size on the unlearning performance during the forgetting process. Because the Pareto-Optimality solver estimates task conflicts and updates the model on a batch basis, the batch size plays a critical role in determining its effectiveness.
Here we choose batch size $\in\{32,64,128,256,512\}$ to conduct experiments. The figure 2 illustrates the performance of the three models on the Games dataset. We clearly observe that as the batch size increases, the U-score tends to decrease. This is because the Pareto-Optimality solver dynamically searches for the optimal gradient update on a per-batch basis. And a smaller batch size reduces conflicts among data samples within each batch, increasing the likelihood of identifying accurate update directions, thereby yielding higher U-scores, but may lead to longer training time. We also measure the training time per epoch under different batch sizes and find that training time decreases as the unlearning batch size grows. For instance, on BERT4Rec, the training time per epoch is 0.4 seconds with a batch size of 256, whereas it is 2.5 seconds with a batch size of 32. The model generally converges to satisfactory unlearning performance after approximately 100 epochs. So this demonstrates a trade-off between unlearning effectiveness and efficiency when selecting the unlearning batch size.
5.5.2. Impact of unlearning data ratio
In this part, we investigate the performance of different models under varying amounts of unlearning data, as illustrated in the figure 3. GRU4Rec is chosen as the backbone recommender model and trained on Beauty dataset. We set the unlearning data ratio of the training data from 3% to 10%. We can observe that the variation in U-score for Retrain is minimal, as it represents the most ideal yet computationally expensive approach. CAU also exhibits a variation within 1% when the ratio changes, further indicating the greater stability and robustness of our method, particularly due to the contributions of the Pareto-Optimality solver and sampling strategy. In contrast, the stability of SISA and SRU decreases as the amount of unlearning data increases, with SISA dropping by 8% and SRU by 4%.
Figure 3. Impact of the unlearning data ratio.
6. Conclusion
In this paper, we propose CAU, a curriculum-based approximate unlearning framework specifically designed for session-based recommendation systems. By formulating unlearning as a multi-objective optimization problem and seeking Pareto-Optimal solutions, CAU effectively balances the trade-off between forgetting targeted samples and preserving overall recommendation performance. Furthermore, the curriculum learning strategy, guided by the proposed gradient and embedding unlearning difficulty scores, along with the hard-sampling and soft-sampling methods, enables a more efficient and effective unlearning process. Extensive experiments on multiple benchmark datasets demonstrate the superiority of CAU over existing methods, confirming its practical value and robustness in handling multiple unlearning requests.
Ethical considerations
Our proposed method for curriculum approximate unlearning in session-based recommendation systems aims to remove the influence of specific user data efficiently. While this can enhance user privacy and control over personal data, potential negative societal impacts exist. For example, improper unlearning could inadvertently degrade recommendation quality for certain users. However, these risks are minimal and can be effectively managed.